I need to find the intercepts for -4x + 10y = 20. I dont know how to solve it. i tried and got yx= 2(not even a possible answer -_-') and so that shows i need help beacuse the answer is suposed to be points.thank you and please hurry.
Answer by Alan3354(69443)

I need to find the intercepts for -4x + 10y = 20.
-----------
The line crosses the y-axis when x = 0.
-4x + 10y = 20.
0 + 10y = 20
y = 2 or (0,2) is the y-int.
----------
The line crosses the x-axis when y = 0.
-4x = 20
x = -5 or (-5,0)